# Architecture for coherent dual-comb spectroscopy and low-noise photonic microwave generation using mechanically actuated soliton microcombs

### v1

@@ Abstract
Microresonator-based optical frequency combs, now known as microcombs, have various applications thanks to their fascinating features including compactness, high-repetition rate, low-power operation, and a broadband spectrum covering up to one-octave [1, 2, 3]. Specifically, the high repetition rates of microcombs ranging from gigahertz to several hundreds of gigahertz have greatly accelerated the acquisition speed of dual-comb spectroscopy [4, 5, 6, 7, 8, 9, 10] and ranging (LiDAR) [11, 12, 13]. Dissipative Kerr soliton (DKS) states provide a mode-locked pulse stream with a femtosecond pulse width alongside a broadband optical spectrum sustained by the double balance of dispersion and nonlinearity as well as parametric gain and cavity loss. Such unique features of DKSs have been utilized in dual-comb technologies to achieve accuracy, a high acquisition rate, and a small footprint simultaneously. To realize dual-DKS combs, we need twin microresonators with slightly different mode spacings, but achieving mutual coherence between two DKS sources is a challenge in this system [4, 5, 6]. As an alternative approach, counter-propagating [14, 15, 16] and spatial multiplexing DKSs [7, 17, 16] have been employed to ensure coherence by unifying a pump laser and a resonator. Although these approaches have several advantages over the twin-resonator configuration, they unfortunately sacrifice flexibility in dual-combs such as that related to frequency differences in repetition rates and carrier offsets that are essential if we are to fully exploit dual-comb techniques [7, 8, 9, 10].
@@
Figure 3(a) shows optical spectra of dual soliton combs, where the solitons share the same pump laser. The repetition rates () of each comb are monitored with a high-speed photoreceiver and an electrical spectrum analyzer (Fig. 3(b) and 3(c)), they are 9.36267 GHz and 9.40149 GHz, respectively. Here, a mechanically actuated comb shows a slightly higher repetition rate, and the difference in repetition rates () corresponds to 38.82 MHz. This result indicates a compression factor of and the measurable bandwidth without spectral aliasing is THz.
@@
Two individual soliton combs are combined by using a 50 50 fiber coupler after they have passed through a fiber-Bragg grating filter to eliminate any residual pump light. An interference optical spectrum is presented in Fig. 4(a). We observe an interferogram spectrum in the electrical domain, i.e., RF comb, as a result of down-conversion from the optical domain (Fig. 4(b) and 4(c)). The rf comb lines are equidistantly separated by 38.82 MHz with a signal-to-noise ratio of 30 dB. A close-up view of one of the RF combs (Fig. 4(d)) indicates a significant narrow linewidth, which is comparable to the beat-note linewidth of the soliton repetition rate. In contrast, the RF comb exhibits a much broader linewidth when distinct pump lasers are used to generate dual combs due to the absence of coherence and relative frequency drift as shown in Fig. 4(e). This is clear evidence that dual combs have mutual coherence without coherence degradation when we use different pump lasers. Thanks to the use of a single-pump, two-cavity configuration, we can enhance the flexibility of microresonator-based dual comb techniques, for instance, the controllability of while maintaining both mutual coherence and system simplicity. In principle, the difference in repetition rates can be precisely controlled by employing dispersion engineering [31] and ultraprecision machining [32] and aligned by thermal tuning with a rate of .
@@
Next, we present a low-noise DKS directly pumped by a USL. Earlier studies have revealed that soliton microcombs receive several noise contributions from including the phase and intensity noise of the pump laser, shot noise, and thermo-refractive noise [25]. Importantly, each comb line of the DKS directly inherits the noise properties of the pump laser owing to coherent parametric processes, and therefore a promising way of reducing the comb noise is to exploit an ultra-stable, ultralow-noise CW laser as the pump laser. Here, we use an ultra-stable reference laser (Menlo Systems, ORS1500) with a wavelength of 1542 nm (Fig. 1), where a CW laser is tightly locked to a high-finesse cavity made of ultra-low expansion (ULE) glass, ensuring an ultimate low-noise pump source with a linewidth of less than 1 Hz. Although the frequency of the USL cannot be tuned, mechanical actuation readily allows the generation of soliton microcombs.
Figure 5(a) shows the observed spectrum of a single soliton state followed by the optical amplification and polarization adjustment of the pump laser (Fig. 2(b)). The inset shows the beat-note spectrum, yielding a repetition rate of 15.24 GHz.
@@
Of particular interest is the phase noise of the free-running soliton repetition rate because the USL is expected to exhibit an outstanding phase noise property compared with those of commonly-used pump sources such as an external cavity diode laser or a fiber laser. Figure 5(b) shows the phase noise power spectrum density (PSD) of the fundamental repetition rate. At offset frequencies of 100 Hz to 100 kHz, the USL-pumped DKS exhibits greatly suppressed phase-noise performance compared with the DKS pumped by a fiber laser (NKT Photonics, Koheras E15). It is noteworthy that the phase noise decreases by up to approximately 50 dB at around 5 kHz, reaching a lower limit imposed by the measurement floor. The PSD has a gentle peak at 100 kHz, where a similar peak can be found in the optical phase noise of the USL due to the servo control bandwidth, i.e., servo bump. This indicates that the soliton microcomb inherits the characteristics of the USL. It should be noted that large variations at low offset frequencies ( Hz) arise from cavity temperature fluctuation, which can be compensated for by applying pump-detuning stabilization; the PSD is clearly reduced by supplying the PDH feedback to the pump laser frequency below its control bandwidth ( a few kHz). In short, the phase noise of a soliton microcomb is significantly suppressed thanks to the superior frequency stability inherent to the USL.

### v2

@@ Abstract
Microresonator-based optical frequency combs, now known as microcombs, have various applications thanks to their fascinating features including compactness, high-repetition rate, low-power operation, and a broadband spectrum covering up to one-octave [1, 2, 3]. Specifically, the high repetition rates of microcombs ranging from gigahertz to several hundreds of gigahertz have greatly accelerated the acquisition speed of dual-comb spectroscopy [4, 5, 6, 7, 8, 9, 10] and ranging [11, 12, 13]. Dissipative Kerr soliton (DKS) states provide a mode-locked pulse stream with a femtosecond pulse width alongside a broadband optical spectrum sustained by the double balance of dispersion and nonlinearity as well as parametric gain and cavity loss. Such unique features of DKSs have been utilized in dual-comb technologies to achieve accuracy, a high acquisition rate, and a small footprint simultaneously. To realize dual-DKS combs, we need twin microresonators with slightly different mode spacings, but achieving mutual coherence between two DKS sources is a challenge in this system [4, 5, 6]. As an alternative approach, counter-propagating [14, 15, 16] and spatial multiplexing DKSs [7, 17, 16] have been employed to ensure coherence by unifying a pump laser and a resonator. Although these approaches have several advantages over the twin-resonator configuration, they unfortunately sacrifice flexibility in dual-combs such as that related to frequency differences in repetition rates and carrier offsets that are essential if we are to fully exploit dual-comb techniques [7, 8, 9, 10].
@@
Figure 2(b) shows optical spectra of dual soliton combs, where the solitons share the same pump laser. The repetition rates () of each comb are monitored with a high-speed photodetector and an electrical spectrum analyzer (Fig. 2(c) and 2(d)), they are 9.36267 GHz and 9.40149 GHz, respectively. Here, a mechanically actuated comb shows a slightly higher repetition rate, and the difference in repetition rates () corresponds to 38.82 MHz. This result indicates a compression factor of and the measurable bandwidth without spectral aliasing is THz.
@@
Two individual soliton combs are combined by using a 50 50 fiber coupler after they have passed through a fiber-Bragg grating filter to eliminate any residual pump light. An interference optical spectrum is presented in Fig. 3(a). We observe an interferogram spectrum in the electrical domain, i.e., RF comb, as a result of down-conversion from the optical domain (Fig. 3(b) and 3(c)). The RF comb lines are equidistantly separated by 38.82 MHz with a signal-to-noise ratio of 30 dB. Here, dual combs share the same pump laser and the comb lines on both sides with respect to the pump frequency will have equal heterodyne signals, making them indistinguishable in the RF domain. To avoid this limitation, a frequency shifter (e.g., an acousto-optic modulator) can be added on one side of dual combs. A close-up view of one of the RF combs (Fig. 3(d)) indicates a significant narrow linewidth, which is comparable to the beat-note linewidth of the soliton repetition rate. In contrast, the RF comb exhibits a much broader linewidth when distinct pump lasers are used to generate dual combs due to the absence of coherence and relative frequency drift as shown in Fig. 3(e). This is clear evidence that the coherence of the dual combs would degenerate when we use different pump lasers. Thanks to the use of a single-pump, two-cavity configuration, we can enhance the flexibility of microresonator-based dual comb techniques, for instance, the controllability of while maintaining both mutual coherence and system simplicity. In principle, the difference in repetition rates can be precisely controlled by employing dispersion engineering [31] and ultraprecision machining [32] and aligned by thermal tuning with a rate of .
@@
Next, we present a low-noise DKS directly pumped by a USL. Earlier studies have revealed that soliton microcombs receive several noise contributions from including the phase and intensity noise of the pump laser, shot noise, and thermo-refractive noise [33, 25]. Importantly, each comb line of the DKS directly inherits the noise properties of the pump laser owing to coherent parametric processes, and therefore a promising way of reducing the comb noise is to exploit an ultra-stable, ultralow-noise CW laser as the pump laser. Figure 4(a) shows the experimental setup. Here, we use an ultra-stable reference laser (Menlo Systems, ORS1500) with a wavelength of 1542 nm, where a CW laser is tightly locked to a high-finesse cavity made of ultra-low expansion (ULE) glass, ensuring an ultimate low-noise pump source with a linewidth of less than 1 Hz. Although the frequency of the USL cannot be tuned, mechanical actuation readily allows the generation of soliton microcombs. Figure 4(b) shows the observed spectrum of a single soliton state followed by the optical amplification and polarization adjustment of the pump laser (Fig. 4(a)). The inset shows the beat-note spectrum, yielding a repetition rate of 15.24 GHz. It should be noted that no active feedback such as detuning stabilization is performed in this experiment.
@@
Of particular interest is the phase noise of the free-running soliton repetition rate because the USL is expected to exhibit an outstanding phase noise property compared with those of commonly-used pump sources such as an external cavity diode laser or a fiber laser. Figure 4(c) shows the phase noise power spectrum density (PSD) of the fundamental repetition rate. At offset frequencies of 100 Hz to 100 kHz, the USL-pumped DKS exhibits greatly suppressed phase-noise performance compared with the DKS pumped by a fiber laser (NKT Photonics, Koheras E15). It is noteworthy that the phase noise decreases by up to approximately 50 dB at around 5 kHz, reaching a lower limit imposed by the measurement floor. The PSD has a gentle peak at 100 kHz, where a similar peak can be found in the optical phase noise of the USL due to the servo control bandwidth, i.e., servo bump. This indicates that the soliton microcomb inherits the characteristics of the USL. It should be noted that large variations at low offset frequencies ( Hz) arise from cavity temperature change and subsequent detuning fluctuation, which can be compensated for by applying pump-detuning stabilization. We confirmed that the variation in the PSD is clearly reduced when the PDH feedback is supplied to the pump laser frequency below its control bandwidth ( a few kHz) as shown by the solid red line in Fig. 4(c). In short, the phase noise of a soliton microcomb is significantly suppressed thanks to the superior frequency stability inherent to the USL. Further improvements in the SSB phase noise can be expected if we operate the DKS at a quiet point [23, 34] with pump-detuning stabilization.